Suppose that A, B, C are 3×3 matrices with det (A) = 2, det (B) = 3 and det (C) = 5. Compute the following determinants:

(a) det (AB)

(b) det (3AB-2C2)

(c) det (A2CTB-1)


1
Expert's answer
2021-09-09T05:01:29-0400

ANSWERS

"(a) \det (AB) = 6 \\
(b) \det (3AB^{-2}C^2) = 150 \\
(c) \det (A^2C^TB^{-1}) = \frac{20}{3}"


SOLUTIONS

"(a) \det(AB) = \det(A) \times \det(B)"

"= 2 \times 3"

"=6"


"(b) \det(3AB^{-2}C^2) = 3^3 \cdot \det(A) \cdot \det(B)^{-2} \cdot \det (C)^2"

"= 27 \times 2 \times \frac{1}{3^2} \times 5^2"

"= 27 \times 2 \times \frac {1}{9} \times 25"

"= 150"


"(c) \det(A^2C^TB^{-1}) = \det(A)^2 \cdot \det(C^T) \cdot \det (B^{-1})"

"= 2^2 \times 5 \times \frac {1}{3}"

"= 4 \times 5 \times \frac {1}{3}"

"= \frac {20}{3}"